Un-polarized light in air is incident upon Lucite (refraction index of 1.50) in such that the reflected light is horizontally polarized. Calculate thefollowing ,00 A a. The incident angle of the light from the surface normal. 27M b. The angle relative to normal of the light entering the Lucite く.so) er
Answer #1

Using Brewster's law,

θ = tan-1(n2)

\large \theta =tan^{-1}(\frac{1.5}{1})

Brewster angle =56.31°

Refracted angle.

\large n1sin i= n2 sin r \\ ?? 1sin 56.31 = 1.5 sin r \\

\large => 1 sin 56.31=1.5 sin r \\ => r= sin^{-1}(0.554700632)

= Refracted angle = 33.69°.

Part A -

From surface normal angle of incidence 56.31° counter clockwise.

Part B-

Angle normal to the light entering lucas is horizontal angle that is reflected light.

thanks .

best wishes